OpenStudy (anonymous):

x+x=2x is this possible if not what the correct way

OpenStudy (turingtest):

x+x=2x is absolutely true and correct always will be, just like 1+1=2

OpenStudy (anonymous):

what property is it

OpenStudy (turingtest):

combining like terms I'll look up the correct name, hold on...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

I don't see this fitting into one if the fundamental properties, see if you can find one that looks like it fits on this site. http://www.capitan.k12.nm.us/teachers/shearerk/basic_rules_of_algebra.htm

OpenStudy (anonymous):

ya i dont either so what do i say it is

OpenStudy (anonymous):

what are 2 examples of this

OpenStudy (turingtest):

BRB sorry...

OpenStudy (turingtest):

Not sure what you mean, but from this follows the ideas x+3x=4x 7y-2y=5y You can prove this with the distributive property: x+3x=x(1+3)=x(4)=4x

OpenStudy (anonymous):

i have 2 more 2 squared =-9 and txt=2x

OpenStudy (anonymous):

x times x =2x

OpenStudy (turingtest):

I'm sorry I have to leave. You should repost your question so someone else can answer. BTW x*x=x^2 not 2x